A raindrop of mass 3x10^-5kg travelling at speed of 20miles/h strikes a walking man of 64kg with velocity -15m/s. Find the final velocity.



1
Expert's answer
2021-07-15T12:39:56-0400
"p_x=(64)(15)=960\frac{kgm}{s}\\p_y=(3\cdot10^{-5})(20)0.447=0.00027\frac{kgm}{s}\\p=960\frac{kgm}{s}\\\theta=\arctan{\frac{0.00027}{960}}=2.8\cdot10^{-7}\degree"